OpenStudy (anonymous):

Confused about complex number arguments. To convert from regular form (a+bi) to exponential form re^iθ, θ=tan^-1(b/a) right? For the complex number -1/2+i(sqrt3)/2, θ=tan^1(-sqrt3)=-pi/3. When inputting e^(-pi/3)i back into my calculator it gives an answer of 1/2-i(sqrt3)/2, the signs are wrong! I worked out that it's the other solution of tan^1(-sqrt3), θ=2pi/3 that gives an answer of -1/2+i(sqrt3)/2 (the right one). My question is, how can I tell which solution of tan^-1(b/a) to use? Thanks.

OpenStudy (anonymous):

Well you're talking about the complex number: \[\Large - \frac{1}{2}+\frac{\sqrt{3}}{2}i\] So if you compute an argument, you want to make sure that you get the precise one, this number lies in the 2nd quadrant, so for that quadrant you have: \[\Large 90° \leq \varphi \leq 180° \]

OpenStudy (anonymous):

Or if you prefer working in radians: \[\Large \frac{\pi}{2} \leq \theta \leq \pi \]

OpenStudy (anonymous):

Ah yes, I forgot about taking the quadrants into account. Thanks!

OpenStudy (anonymous):

you're welcome